Consider the utility function π‘ˆ(π‘₯, 𝑦) = π‘₯𝑦 + π‘₯, with π‘€π‘ˆπ‘₯ = 𝑦 + 1 and π‘€π‘ˆπ‘¦ = π‘₯. (Verify these.)Β 

Β 

a. is the assumption of β€œmore is better” (non-satiation) satisfied for both goods x and y?Β 

b. does the marginal utility of π‘₯ diminish, remain constant, or increase as the consumer buys more π‘₯? Explain.Β 

c. what is 𝑀𝑅𝑆π‘₯,𝑦?Β 

d. is 𝑀𝑅𝑆π‘₯,𝑦 diminishing, constant, or increasing as the consumer substitutes π‘₯ for 𝑦 along an indifference curve?Β 

e. on a graph with π‘₯ on the horizontal axis and 𝑦 on the vertical axis, draw a typical indifference curve (not necessarily exactly to scale, but should reflect whether 𝑀𝑅𝑆π‘₯,𝑦 is diminishing or not). Label the curve as π‘ˆ1.Β 

f. on the same graph, draw a second indifference curve π‘ˆ2, where π‘ˆ2 > π‘ˆ1


1
Expert's answer
2021-10-05T17:53:44-0400

Given the utility functionΒ "U(x, y)=xy+x"

"MU_x=\frac{dU}{dx}=\frac{d}{dx}(xy+x)=y+1\\
MU_x=\frac{dU}{dy}=\frac{d}{dy}(xy+x)=x\\ 
Hence \space verified."

a)

Β Since the utility function is increasing in both goods X and Y, hence it can be concluded that the assumption "more is better" is satisfied for both goods X and Y.

b)

"MU_x=y+1",Β is constant with respect to x.

Hence the marginal utility of x remain constant as the consumer buys more x.

c)

"MRS_{x,y}=\frac{\frac{dU}{dx}}{\frac{dU}{dy}}\\\space \\
MRS_{x,y}=\frac{y+1}{x}"

d)

the MRS is decreasing as consumer substitutes X for Y

e)



f)